Elementary Number Theory

Show that if n is prime, then Ø(n) +o(n) = nt(n).
Answer #1

3 show that if n is brime, then din) to (n=nTon let n=b L.H.S =) Acn)toen) = (-1)+ 0() =(p-1) + OC b). = (p-1)+($+_1) (6-1) 1